OpenStudy (anonymous):

\[\int\limits_{}^{} (6x-11 / x+2) dx\]

OpenStudy (amistre64):

nice job with the equation writing skill...kudos ;)

OpenStudy (amistre64):

frac{top}{bottom} makes a clean fraction

OpenStudy (anonymous):

You need to substitute x for x=u+2 and then solve (dx=du since d/dy of +2 is zero).

OpenStudy (anonymous):

Err... that's not d/dy of +2 but the derivative of +2

OpenStudy (amistre64):

\[\frac{top}{australia}\]

OpenStudy (amistre64):

\[\int \frac{6x-11}{x+2}dx\]

OpenStudy (anonymous):

thanks for the lesson in writing equations >:(

OpenStudy (anonymous):

Do you get what I wrote msunprecedented?

OpenStudy (anonymous):

no, i did not

OpenStudy (amistre64):

lol.... I was impressed by your talent ...still am :)

OpenStudy (amistre64):

do you want the integral of 3 seperate terms; or is what I posted a correct interpretation of it?

OpenStudy (anonymous):

no its correct \[\int\limits_{}^{} \frac {6x-11} {x+2} dx\]

OpenStudy (amistre64):

... factoring might be possible but we would have to find out what value had been canceled before hand.... and then do partial decomposition as a possibility

OpenStudy (anonymous):

Actually, this is really simple if you view the fraction like this: \[6x/(x+2) - (11/x+2)\] Now that it's in this form, you split the integral into two : \[\int\limits_{?}^{?}6x/(x+2)dx - \int\limits_{?}^{?}11/(6x+2)dx\] This should be easier to integrate... let me know if you need more help

OpenStudy (amistre64):

the left is easy now :) the right still has a straggler.... that "x" up top can get bothersome

OpenStudy (amistre64):

and by left I mean right ;)

OpenStudy (amistre64):

would the uv substitution work well there?

OpenStudy (anonymous):

uv substitution wouldn't work because I don't think you could factor x+2. (I can't see a way to factor anyway)

OpenStudy (amistre64):

6x goes to zero pretty quick like; but I get ln(x+2) and get lost on integrateing that

OpenStudy (anonymous):

Hmm.. I think integration by substitution works here. If you let u=x+2, you can get the right equal to:\[\int\limits_{?}^{?} (6u-12)/u *du\] Then you split the fraction one more time

OpenStudy (anonymous):

I mean the left when I say the "right" above

OpenStudy (amistre64):

lol..... stage left ...STAGE left!! :)

OpenStudy (anonymous):

Yikes! I'm soo confused! U-sub here? Integration by parts there? :(

OpenStudy (anonymous):

msunprecedented, have you learnt how to do integration by substitution yet?

OpenStudy (anonymous):

I believe not.